hartnn (hartnn):

If a+ib=tanh(x+i π/4),prove that a^2+b^2=1

hartnn (hartnn):

what is the shortest way....taking log on both sides? didn't help much...

OpenStudy (anonymous):

first expand the right side to the definition of tanh

OpenStudy (anonymous):

i would do it the difficult way :P

hartnn (hartnn):

\(\Large \dfrac{e^y-e^{-y}}{e^y+e^{-y}}\) where y =x+i π/4 then ?

hartnn (hartnn):

\(\Large e^y = e^{x+i\pi/4} = e^x e^{i\pi/4} = ...\)

OpenStudy (anonymous):

well i would probably condense it to -i/cot(i*y)

hartnn (hartnn):

oh and then ?

hartnn (hartnn):

any formula for cot iy ?

OpenStudy (anonymous):

so we have a + bi = -i/cot(i*y) multiply both sides by i ai - b = 1/cot(i*y) ai - b = tan(i*y) ai - b = tan(ix - pi/4)

OpenStudy (anonymous):

what do you think you can do from there?

hartnn (hartnn):

tan (a-b) formula ? but that will make the denominator have the complex term...

OpenStudy (anonymous):

yeah, and there's an i on the left side of the equation as well. we can multiply things out to make it work (i think, i may have done it wrong but i think it'll work :P trust)

hartnn (hartnn):

sure! lets try ai-b = (tan ix-1)/(1-tanix) = -1 whoa...is that correct ?

OpenStudy (anonymous):

yes! (tan (ix)-1)/(1-tan(ix)) = -1

hartnn (hartnn):

thanks!

OpenStudy (anonymous):

no worries. you can prove it with the given info right?

hartnn (hartnn):

ofcourse :)

hartnn (hartnn):

the first step was important!

OpenStudy (anonymous):

yeah, i hate working with tanh identities. much easier to just expand tanh because -i/cot(ix) is a very pretty way of putting tanh